OpenStudy (anonymous):

WILL GIVE MEDAL Write an equation with a variable on both sides of the equal sign that has infinitely many solutions. Solve the equation and explain why it has an infinite number of solutions.

OpenStudy (anonymous):

x-4*3=x-12 This would be just one example of infinitely many solutions, because in order to have infinitely many solutions, you need to have both sides of the equation be the same, the question requires, though, that you have a variable on both sides, you can use this equation because it has a variable on both sides, and when simplified, looks like this... x-12=x-12 Both sides are the same and so have infinitely many solutions.